Using revtex.bst with REVTEX 4

October 11, 2000

Contents

1 Introduction

With the introduction of REVTEX 4, there is an emphasis on tagging information in manuscripts so that it can be reliably extracted and used in the editorial and production processes. One of the primary areas where such tagging can have a major impact in is in the bibliography sections of manuscripts. With reliable tagging, references can be automatically checked for accuracy and used for creating hyperlinks to cited papers available online. Improved accuracy in references will filter all the way through the production process to the online journal products. Production costs and delays can be greatly reduced by catching errors at the earliest stage possible, ideally at the time of submission. Because tagging information can be tedious, we have created a BibTEX style file that handles all of the tagging for the author. One only has to supply a proper .bib file and use the revtex.bst style file.

Readers unfamiliar with BibTEX may wish to consult Sec. ?? before reading further.

1.1 Reference component tagging

The components that build up a reference are each tagged using a single command, \bibinfo. A component corresponds to a field in an entry in the .bib file. The first argument of \bibinfo is a label that for the most part corresponds to the field names in the .bib entries. The second argument is the value of the field in the .bib entry. For example, if an article entry has its journal set to “Phys. Rev. B”, revtex.bst will output

   \bibinfo{journal}{Phys. Rev. B}.

Usually the \bibinfo command has no effect upon the typesetting. It is only there to provide tagging for later use. However, specific fields can have behaviors associated with them. For instance, the URL field can be typeset using \texttt or even turned into a hyperlink when one uses the hyperref package.

1.2 Differences with prsty.bib

revtex.bst is a complete rewrite of prsty.bib. For ease of maintenance, the new style file was created using Patrick Daly’s custom-bib BibTEX style generating package. The final output forms of the entries will be the same as with prsty.bib, although particular fields may be typeset slightly differently (as with URLs).

1.2.1 New fields

New fields have also been added to revtex.bst. These fields can appear in any of the entry types in Sec. ??.

• collaboration: A collaboration name associated with a reference.

• eprint: The e-print citation for the reference (Los Alamos or otherwise). E-print numbers can be turned into hyperlinks to the online article.

• URL: A URL associated with the reference. Can be turned into a hyperlink.

A collaboration appears as part of the author list and the other two appear at the end of the citaton.

2 Getting started with BibTEX

Standard L^{A}TEX 2εdistributions should include everything that you need to use BibTEX. Here we give a brief overview of BibTEX basics to get you oriented. More information on BibTEX can be found in Leslie Lamport’s book, LaTeX, A Document Preparation System and A Guide to L^{A}TEX 2ε by Helmut Kopka and Patrick W. Daly.

2.1 Files associated with BibTEX

BibTEXhas two main files that an author needs to be familiar with. The first is the .bib file which contains a database, or list, of entries for each item you plan to cite in your paper. Examples of entries are given in ??. The second is a .bbl file which is created by running BibTEX on your manuscript. The .bbl file actually contains the TEX that will be read into your manuscript at the appropriate place.

2.2 Citing BibTEXentries

Each entry in your .bib file has a unique label associated with it. To cite a reference, use the \cite command with the label as an argument. For example, if an @Article entry has the label “impurity,” use the command \cite{impurity}. BibTEX and L^{A}TEX 2ε will take care of numbering and ordering the citations (based partially upon the information in the revtex.bst style file).

2.3 Referencing the .bst and .bib files

To use the revtex.bst style file, put the command \bibliographystyle{revtex} after the \begin{document} line of the manuscript. At the point in the manuscript where the bibliography is to appear, put the command \bibliography{filename} where filename is the basename of your .bib file (i. e., the filename without the .bib extension).

2.4 Running BibTEX

The first step is to create an .aux file by running L^{A}TEX 2ε on your TEXfile. Among other things, this outputs a list of the labels that used in manuscript’s \cite commands. BibTEX uses the .aux file to pull the correct entries out of the .bib file, processes them according to the .bst file, and then puts the results in the .bbl file. After running L^{A}TEX 2ε once, run BibTEX (for example, by typing bibtex filename, where filename is the name of your manuscript’s TEX file without the .tex extension). Once the .bbl file has been created, run L^{A}TEX 2ε again as usual until the warnings about re-running to get the cross-references correct disappear. The .bbl file only needs to be re-created when a reference is added or deleted from the manuscript.

2.5 Entry types

All of the standard entry types are supported. Here we give some example entries that one might put in a .bib file. Please note that some of the data is completely fictitious and these are just given as examples. For more examples, please see more definitive BibTEX references. The eprint and url fields can appear in any of the entry types.

2.5.1 Article

@Article{Distler:1992gc,
     author    = "Jacques Distler and Mark D. Doyle",
     title     = "World sheet supersymmetry without contact terms",
     journal   = "Phys. Lett.",
     volume    = "B287",
     year      = "1992",
     pages     = "73-81",
     eprint    = "hep-th/9203058",
 }
@Article{Mart1984,
     author = "J. Martinsen and S. M. Palmer and J. Tanaka and R. C. Greene
                  and B. M. Hoffman",
     title = "",
     journal = "Phys. Rev. B",
     volume = "30",
     pages = "6269",
     year = "1984",
}

2.5.2 Books, proceedings, and things therein

@Book{Abrik1972,
        author = "A. Abrikosov",
        title = "Introduction to the Theory of Normal Metals",
        publisher = "Academic Press",
        address = "New York",
        year = "1972",
}

2.5.3 Unpublished

@Unpublished{SeiWit1999,
     author    = "Nathan Seiberg and Edward Witten",
     title     = "The D1/D5 System And Singular CFT",
     year      = "1999",
     eprint    = "hep-th/9903224",
 }

2.5.4 Theses

@PHDThesis{Thomp1995,
        author = "J. Thompson",
        title  = "Heavy Metal",
        school = "Northwestern University",
        year = "1995",
        eprint = "math-ph/9805004",
}

2.5.5 Miscellaneous

@Misc{revtex,
      author = "The American Physical Society",
      title  = "The REVTeX 4 home page",
      year   = "1999",
      url    = "http://publish.aps.org/revtex4/",
}